Design a 5-tap FIR bandpass filter with a lower cutoff frequency of
1,600 Hz, an upper cutoff frequency of 1,800 Hz, and a sampling rate
of 8,000 Hz using

a. rectangular window function

b. Hamming window function.

Determine the transfer function and difference equation of the designed
FIR system, and compute and plot the magnitude frequency response
for Ω= 0, π/4, π/2, 3π/4, and π radians.

PLEASE SHOW STEPS CLEARLY
Answer #1

* a) Denýn of FIR filters using Retangular window :- The desired frequency response of a fitter is given by , Haleiw) {! forha(n)= sin (1) Cochem.-0 in wp-?! Kunrating to N sampler hin) = 4. min () [N = 5] Given Tin Rectangle window, -240 12 for imi- (N-1)/2 Relizable filter, multiply by Z N-5 - 2 > Z M -2 -1 H(Z) = 0.52² » Ž (5-1)/2 #la). *[+ +(0.3181)(2+2] +0.3183 zi 7* Plot the Response: يا : -2 نما indey (0) * = 15° | R2 = 40 an/ 1013E ۔ (مائع) + 4 N=1&2 = iss | 0،049 |-0.34 0.9 گ,0 in d*b). Design of FIR filters wings Hemming Window :- The desired frequency response of a low pass filter is given by, Hd (ew) -; #3: ha(n) = sin / (3 (n-3)] * (n-3) of n=3. ha(3)= j(3-3) 27 stly 1 ہے - جمال -انے stly lidw at dw 2 ves A(n-3) na 3. herecoefficiento M haln) 0.075 Whamins hin) = hain). Wham (n) 0:08 0.006 0 11-0:159/0.54 Quyla -0.048586 2 0.225 0.08 0.018 مر 0.n=3; ha (3) 3/4 0.75 Ww cm (3) 7 0.54 – 9.46 cos [ 20 (3) 0,54, z h (B) = hd (3). Wham (3) -0.75X0.54 201405 n=u; hd (u) = sesaw [H(a) + 2H16) Con aw +2661.0 w] estaw [Co.018) +2 (0.006) 6 w +2.6-0.0858) a -j2u Cosw] 0.018 +0.012 Cos2w of 0.1716 los08- of- 09- as- 180 135 १० Sh (M (or) + APAnswer question